### v3

# Redefining the dielectric response of nanoconfined liquids: insights from water

Abstract
Recent experiments show that the relative dielectric constant $\epsilon$ of water confined to a film of nanometric thickness reaches a strikingly low value of 2.1, barely above the bulk’s 1.8 value for the purely electronic response. We argue that $\epsilon$ is not a well-defined measure for dielectric properties at sub-nanometer scales due to the ambiguous definition of confinement width. Instead we propose the 2D polarizability $\alpha_{\perp}$ as the appropriate, well-defined response function whose magnitude can be directly obtained from both measurements and computations. Once the appropriate description is used, understanding the interplay between electronic and ionic contributions becomes critical, contrary to what is widely assumed. This highlights the importance of electronic degrees of freedom in interpreting the dielectric response of polar fluids under nanoconfinement conditions, as revealed by molecular dynamics simulations.
The relative dielectric constant $\epsilon_{\perp}$ across a nanometric thin film of water was recently measured to be as low as 2.1 for a width of $w\lesssim 1$ nm. This paper by Fumagalli et al. Fumagalli et al. (2018) attracted enormous attention for two main reasons: ($i$) it is a beautiful, impressive tour-de-force experiment, and ($ii$) the mentioned value is strikingly low, barely above the water bulk’s purely electronic response Hill (1963) $\epsilon^{b}_{\infty}=1.8$. The first point is undisputed. It is the second one we want to qualify in this article.
A small value for $\epsilon_{\perp}$ is not a surprising result. Numerous simulations have already shown that values in the $\epsilon_{\perp}<10$ range are to be expected for confined water geometries Zhang et al. (2013); Zubeltzu et al. (2016); De Luca et al. (2016); Schlaich et al. (2016); Zhang (2018). The explanation of the origin of such reduced value appearing in several works Monet et al. (2021); Dufils et al. (2024); Deißenbeck and Wippermann (2023) after that of Fumagalli et al. Fumagalli et al. (2018), including high-quality calculations (see e.g. Ref. Dufils et al., 2024), agree, in essence, with the earlier literature, all based on a combination of a more rigid structure of the water layer close to the interface Zhang et al. (2013); Zubeltzu et al. (2016); De Luca et al. (2016); Schlaich et al. (2016), and the dead-layer effect (capacitors in series) for the thickness dependence Zhang (2018). Alternatively, in a recent work, it was proposed that the dielectric permittivity reduction of water is not connected to the structural alignment of interfacial water molecules, and instead, it emerges from anisotropic long range dipole correlations next to surfaces Olivieri et al. (2021). The reduction of the dielectric permittivity in other H-bonded and non H-bonded solvents was discussed recently Motevaselian and Aluru (2020), highlighting the generality of the observed reduction of the dielectric permittivity under confinement conditions.
The impact of interfaces on the permittivity tensor of water has attracted significant attention, which has stimulated the use of the dielectric tensor as a local function explicitly quantifying its dependence on the distance to the interface using statistical mechanics fluctuation formuli Ballenegger and Hansen (2005); Bonthuis et al. (2011); Olivieri et al. (2021); Deißenbeck and Wippermann (2023). These studies already showed that major changes in the dielectric response of water appear at distances <1<1 nm from a surface, demonstrating the anisotropy of the permittivity tensor, the relevance of the surface-water interactions in determining the permittivity, and the importance of interfacial layers in defining the dielectric response of interfacial water. Despite the intrinsic interest of calculating the spatial dependence of the dielectric permittivity profile, its interpretation in mesoscopic and molecular scales is not straightforward, as it requires an averaging process over molecular layers Olivieri et al. (2021); Deißenbeck and Wippermann (2023); Dufils et al. (2024). Furthermore, while most studies focused on the modification of molecular degrees of freedom such as dipole-dipole correlations or water orientation at surfaces, the impact of confinement on electronic degrees of freedom has escaped scrutiny. One aspect of particular importance, which we address in our work, is whether confinement influences the electronic polarisability of confined water relative to bulk conditions. This is an important question, especially considering the very low permittivity of confined water inferred from experiments and theory. The importance of confinement effects and of molecular degrees of freedom has been discussed more widely in the investigation of fluid flow at the nanoscale. Advances in this area are promising, particularly in the development of novel devices with enhanced properties, which rely to some extent on the breakdown of the macroscopic theories that assume a continuum framework Boya et al. (2024). We argue below that using macroscopically defined properties is also problematic when considering the electrostatic response of fluids under nanoconfinement conditions.
Our purpose here is not improving on the determination of $\epsilon_{\perp}$ but instead stating that it is an ill-defined quantity in the sub-nm regime, and that insisting on its quantification is misleading the discussion of the key variables determining the electrostatic screening of confined fluids. As noted above, a problem emerges when assigning a dielectric permittivity to a molecular layer. This calculation requires the definition of a layer thickness to integrate the local perpendicular dielectric permittivity. The minima of the water center of mass density profiles normal to the surface are often used to define such thickness. As discussed before Fumagalli et al. (2018); Dufils et al. (2024), the actual value of the computed or measured $\epsilon_{\perp}$ is rather arbitrary, as the result depends on establishing an effective film width $w$ that cannot be uniquely defined. Hence, the definition of the film width becomes essentially the answer to an ill-posed question, which may have sensible answers but not observable ones. Think, e.g. asking what is the width of an isolated graphene sheet. In essence, this is a problem arising from carrying a macroscopic-theory language over to the sub-nm regime.
Figure 1: (a) Schematic view of the simulation box, perpendicular to the confining plane. The gray-shaded regions represent the location of the confining soft potentials. Oxygen and hydrogen atoms in the water molecules are represented as red and white spheres, respectively. (b) Oxygen density profiles along the confining direction for 2D densities $\sigma=$ 0.177 Å-2 (black) and 0.195 Å-2 (blue).
The dielectric constant, relating the polarisation $P$ to the applied electric field, quantifies the response of a material within the macroscopic framework of Maxwell’s equations. It is $P$ that is ill-defined due to the ambiguity in the width, as $P$ is given by the dipole moment per unit volume. Instead, the corresponding characterization for a two-dimensional (2D) system is given by the dipole moment per unit area, $\mathcal{P}_{2D}$. We propose that for sub-nm thickness, the 2D description is the correct one to use, the cross-over from the 3D description being determined by the ambiguity in $\epsilon_{\perp}$ becoming comparable with the value itself. The transverse dielectric response is then characterized by the 2D transverse polarizability
$\alpha_{\perp}\equiv\frac{1}{\epsilon_{0}}\frac{\partial\mathcal{P}_{\mathrm{2D}}}{\partial\mathcal{E}^{\mathrm{ext}}_{\perp}}$ (1)
as already proposed for 2D materials Tian et al. (2020), instead of the conventional susceptibility in 3D. As done for the polarizability of molecules in the gas phase, referring to the externally applied electric field $\mathcal{E}^{\mathrm{ext}}_{\perp}$ is the obvious choice here since it makes no sense to refer to the perpendicular field inside a 2D system. For solid-state thin film systems, a “macroscopically averaged” electrostatic potential can be defined as a function of the perpendicular component Colombo et al. (1991), which is routinely used in interface and heterostructure physics. It has also been used in soft matter thin films (e.g. soap/water interfaces Bresme and Artacho (2010)), but the difficulty remains when averaging the field over an ill-defined width.
$\epsilon_{0}\mathcal{E}^{\mathrm{ext}}_{\perp}$ in Eq. (1) equals to the electric displacement field $D_{\perp}$ to be used Zhang and Sprik (2016) when $P$ is unambiguously defined. The defined $\alpha_{\perp}$ has units of length and is analogous to the conventionally defined molecular polarisability, which has units of volume. It can be called the dielectric thickness of the film. It is important to remember key differences when referring to the external field instead of the internal. For that, the definition in Eq. (1) can be extended to 3D as $\epsilon_{0}\alpha_{\mathrm{3D}}\equiv\partial P/\partial\mathcal{E}^{\mathrm{ext}}$, using the conventional 3D polarisation $P$. Both polarisabilities are related by $\alpha_{\mathrm{3D}}=\alpha_{\perp}/w$ in the context of this work  [^1] Both $\alpha_{\mathrm{2D}}$ and $\alpha_{\mathrm{3D}}$ are rank-two tensors but we are considering one element of the diagonal of the former ($\alpha_{\perp}$), and the latter is a number times the identity tensor for 3D bulk water., since $P=\mathcal{P}_{\mathrm{2D}}/w$. $\alpha_{\mathrm{3D}}$ relates to the conventional relative dielectric constant as $\alpha_{\mathrm{3D}}=1-\epsilon^{-1}$, which means that the range of 1 to $\infty$ for $\epsilon$ (from no response to complete screening of the internal field as in an ideal metal) becomes 0 to 1 for $\alpha_{\mathrm{3D}}$.
To illustrate the main ideas conveyed here, we present two sets of simulations for a bilayer liquid film of water confined along the z direction of the simulation box between two soft Lennard-Jones 9-3 separated by 8 Å between their respective origins (see Figure 1). The parameters of the confining potential are set to mimic the interaction of water with solid paraffinKumar et al. (2005). This same system has been extensively studied in previous works Kumar et al. (2005); Zubeltzu et al. (2016); Zangi (2004); Calero and Franzese (2020); Leoni et al. (2021); Han et al. (2010). The bilayer film thickness is at the lower end of the systems measured in Ref. Fumagalli et al., 2018. For classical trajectories obtained using the empirical force field TIP4P/2005 Abascal and Vega (2005) at $T=300$ K under various values of an applied external field, we calculate $\alpha_{\perp}$ from the $\mathcal{P}_{2D}$ obtained from ($i$) the point charges of the TIP4P/2005 model and $(ii)$ from the perpendicular dipole obtained with density-functional theory (DFT) on a sample of snapshots of the classical trajectory, under the same external electrostatic field values. The TIP4P/2005 trajectories are generated within the NVT ensemble using the Nosé-Hoover thermostat involving 10 ns of equilibration time followed by 10 ns of production time. The simulation box, with dimensions 34.39 $\times$34.39 $\times$23.00 Å, contains between 210 and 236 water molecules. This corresponds to a 2D molecular density range of 0.177 Å-2 - 0.200 Å-2, where water is found to be in the liquid phase  [^2] The 2D molecular density range of 0.177 Å-2 - 0.200 Å-2 corresponds to a 3D mass density range of 1.04 g/cm3 - 1.17 g/cm3 in Ref.Zubeltzu et al. (2016). The periodic image interactions along the confining direction are effectively disabled by using the slab command in LAMMPS Thompson et al. (2022), with a volfactor of 3.0. The DFT calculations were done with the Siesta method Soler et al. (2002) using the PBE density functional Perdew et al. (1996). We use the TZP basis set for water described in Corsetti et al. (2013). To assess the accuracy of our simulation setup we determine the polarisability of a water monomer. We obtain a polarisability of 1.6 Å3 in Gaussian units (see Supplemental Material for calculation details), which is very close to the previously reported value of 1.59 Å3 calculated using the same functional Ge and Lu (2017). This result is also in reasonable agreement with the experimentally reported value of 1.45 Å3 Haynes (2016). The remaining technical details for both kinds of calculations, classical and DFT, as well as a detailed description of the structural and other properties of the film as obtained in the simulations can be found in Ref. Zubeltzu et al., 2016.
Figure 2: 2D polarisation $\mathcal{P}_{\mathrm{2D}}$ vs external traverse electric field $\mathcal{E}^{\mathrm{ext}}_{\perp}$ for TIP4P-2005 water classical trajectories (black) and for DFT calculations using the classical TIP4P-2005 under the same external field (red). In blue, we show the pure electronic response obtained from DFT as explained in the text. Inset: 2D density dependence of $\alpha_{\perp}$ obtained with the TIP4P-2005 classical simulations.
Table 1: Two-dimensional normal polarizability of the water film in Å. The experimental value $\alpha_{\perp}^{\mathrm{exp}}$ is obtained from the values of $\epsilon_{\perp}$ and $w$ reported in Ref. Fumagalli et al., 2018. TIP4P/2005 values ($\alpha_{\perp}^{\mathrm{TIP}}$) were obtained from TIP4P/2005 trajectories under various values of the normal electric field. DFT results correspond to calculations under the same field values on a sample of snapshots of the corresponding TIP4P/2005 classical trajectories ($\alpha_{\perp}^{\mathrm{DFT}}$), or snapshots of classical trajectories for TIP4P/2005 under zero field conditions ($\alpha_{\perp\infty}^{\mathrm{DFT}}$). $\left(\alpha_{\perp\infty}/\alpha_{\perp}\right)^{\mathrm{exp}}_{\mathrm{bulk}}$ corresponds to the ratio of polarisabilities as obtained from $\epsilon=78$ and $\epsilon_{\infty}=1.8$ for bulk water.
$\alpha_{\perp}^{\mathrm{exp}}$ $\alpha_{\perp}^{\mathrm{TIP4P}}$ $\alpha_{\perp}^{\mathrm{DFT}}$ $\alpha_{\perp\infty}^{\mathrm{DFT}}$ $\left(\alpha_{\perp\infty}/\alpha_{\perp}\right)^{\mathrm{DFT}}_{\mathrm{film}}$ $\left(\alpha_{\perp\infty}/\alpha_{\perp}\right)^{\mathrm{exp}}_{\mathrm{bulk}}$
4.57.94.5-7.9 5.7 6.1 2.3 0.37 0.45
Figure 2 shows $\mathcal{P}_{2D}$ versus $\mathcal{E}_{\perp}^{\mathrm{ext}}$ for a 2D molecular density of $\sigma$ = 0.177 Å-2, for both TIP4P-2005 and DFT calculations. Regarding the orientation of the molecular dipole moment predicted by the TIP4P/2005 model along the confining direction, two preferential orientations are observed in each water layer (see Supplemental Material). This agrees with the structure reported in previous calculations using first-principles accuracy neural network potentials Dufils et al. (2024). The 2D polarisation for the DFT calculations is slightly larger, consistent with the deformation of the electronic cloud. The resulting values of $\alpha_{\perp}$ are shown in Table 1, and compared with the results of experiment Fumagalli et al. (2018). For the latter, there is an ambiguity since the actual values of capacitance, $\mathcal{P}_{2D}$, or $\alpha_{\perp}$ were not reported, but $\epsilon_{\perp}=2.1$ was given for a range of water film thickness between 8.5 and 15 Å.
The $\alpha_{\perp}$ interval appearing in the Table corresponds to that range of film thicknesses. Both theoretical values of $\alpha_{\perp}$ (TIP4P/2005, DFT) fall within the range inferred from experiments ($\alpha_{\perp}^{\mathrm{exp}}$). The inset of Figure 2 also shows the dependence of $\alpha_{\perp}$ on the 2D density of water molecules obtained from the analysis of TIP4P/2005 classical trajectories. These results address the experimental difficulty in ascertaining the experimental pressure conditions Fumagalli et al. (2018). Our results (see Fig. 2) show that $\alpha_{\perp}$ features a very small dependence on the 2D density.
In addition to directly calculable, the proposed 2D polarisability $\alpha_{\perp}$ can be directly obtained from experiment, in particular from capacitance measurements using $l-\alpha_{\perp}=\epsilon_{0}/C$, where $C$ is the capacitance per unit area, and $l$ is the effective distance between the capacitor plates. This distance is in principle ill-defined at the nanoscale, as the effective width of a plate capacitor depends on the charge distribution around its surface planesManino et al. (2024). Differential capacitance measurements can be used to avoid the dependence on $l$, as done in Ref. Fumagalli et al., 2018. In particular, by measuring the capacitance per unit area with and without dielectric (water), $C$ and $C_{0}$, respectively,
$\alpha_{\perp}=\epsilon_{0}(C_{0}^{-1}-C^{-1}).$ (2)
If a change in the inter-plate distance $\Delta l$ is measured when the capacitor is filled with water (as in Ref. Fumagalli et al., 2018), it simply needs to be added to Eq. 2: $\alpha_{\perp}=\Delta l+\epsilon_{0}(C_{0}^{-1}-C^{-1})$. Note that $\Delta l$ is independent of the choice of the reference position of each plate, as it directly quantifies the displacement that occurs between the plates before and after the filling with water. Additionally, the measurements of $C$ and $C_{0}$ are enough to define an “effective" dielectric constant, $\epsilon^{eff}_{\perp}=\frac{C}{C_{0}}$, without the need to estimate the effective distance $l$. $\epsilon^{eff}_{\perp}$ represents the overall dielectric behavior of the plate capacitor’s interiorBecker et al. (2023); Loche et al. (2020), and it is not the dielectric constant of the water film. It can be used to partition the system into layers and evaluate the total capacitance as a model of capacitors in series. But such partitions, particularly the surface ones are still ill-defined and arbitrary. We illustrate here the extreme sensitivity of $\epsilon_{\perp}$ if insisting on defining a water film width $w$, in spite of the acknowledged arbitrariness of such a choice Dufils et al. (2024). The macroscopic formalism is thereby recovered, with $\alpha_{\perp}$ relating to $\epsilon_{\perp}$ through Zhang and Sprik (2016)
$\epsilon_{\perp}=\left(1-\frac{\alpha_{\perp}}{w}\right)^{-1}\;.$ (3)
$\epsilon_{\perp}$ diverges when $w=\alpha_{\perp}$, which nicely points to the physical meaning of the length $\alpha_{\perp}$: the width of an ideal metal film that would respond with the same polarisability as the one being measured. The divergence is also responsible for the high sensitivity of the value of $\epsilon_{\perp}$ to the chosen $w$. Fig. 3 illustrates that variability for the DFT simulations described above, showing that a given value of $\alpha_{\perp}$ offers dramatically different values of $\epsilon_{\perp}$ for different sensible choices of $w$.
Figure 3: Dielectric constant across a water thin film as a function of the defined width $w$ of the film for a given value of the 2D polarisability, $\alpha_{\perp}=6.06$ Å, as obtained with DFT simulations described in the text. Values of $\epsilon_{\perp}$ are highlighted for three values of $w$ for the same film: (ii) the distance between the origins of the confining potentials $w$ = 8 Å, (iiii) $w=5.17$ Å, the accessible perpendicular space as defined by Kumar et al. Kumar et al. (2005), for this bilayer film, and (iiiiii) $w=5.31$ Å, obtained from normalising the 2D density of the film with the 3D density of bulk water. Negative values for $\epsilon_{\perp}$ indicate overscreening, a larger polarisation than what is needed to screen the field completely.
The 2D polarisability $\alpha_{\perp}$ as defined here is an intensive property in the plane (per unit area), but extensive out of plane, and it depends on the amount of substance per unit area and the 2D density $\sigma$. When it is normalized by the experimental density $\rho$ of bulk water at ambient conditions, it gives another effective thickness, $w=\sigma/\rho$, which converges to the film thickness in the macroscopic limit.
In our simulations, we obtain $w=5.31$ Å for the molecular 2D density $\sigma$ = 0.177 Å-2 and $T=300$ K. We include this thickness in the analysis shown in Fig. 3.
Finally, we address the striking experimental result that the reported dielectric response of the film $\epsilon_{\perp}=2.1$ is only marginally larger than the electronic dielectric constant of bulk water $\epsilon^{b}_{\infty}=1.8$.
We calculate the electronic 2D polarizability $\alpha_{\perp\infty}^{\mathrm{DFT}}$ of the film by running DFT calculations under different values of $\mathcal{E}_{\perp}^{\mathrm{ext}}$ on a sample of snapshots from the zero-field TIP4P/2005 trajectory. The results are shown in Fig. 2 giving the value of $\alpha_{\perp\infty}^{\mathrm{DFT}}=2.3$ Å (Table 1), which corresponds to a ratio of $\alpha_{\perp\infty}^{\mathrm{DFT}}/\alpha_{\perp}^{\mathrm{DFT}}$ of 0.37 for the film. This ratio represents the percentage of electronic to total out of plane polarisability in the film. It can also be estimated for bulk water from experiments as:
$\left(\alpha^{\mathrm{3D}}_{\infty}/\alpha^{\mathrm{3D}}\right)^{\mathrm{exp}}_{\mathrm{bulk}}=(1-\epsilon^{-1}_{\infty})/(1-\epsilon^{-1}).$ (4)
[…]
The indications obtained in this work point to a reduction of the electronic response of the film. While this provides a plausible explanation for the striking experimental result reported, further investigation would be desirable to confirm and quantify this effect more precisely. In particular, ab initio molecular dynamics simulations could offer valuable insight. In this sense, our observations agree with the comment of Ref. Olivieri et al. (2021) that the reduction in $\epsilon_{\perp}$ is not necessarily connected to interfacial water alignment. Our results suggest instead that it may be related to changes in electronic polarizability, a point that has not been considered or addressed in previous works.
That reduction becomes the intriguing result now. The directional dependence of the electronic polarizability tensor of the water molecule Ge and Lu (2017) is too small to support the hypothesis of the electronic response reduction being due to prevalent molecular orientations in the film.
Nonetheless, the thickness dependence observed for $\epsilon_{\perp}$ remains nicely explained by the capacitors-in-series, dead-layer picture. The point that it is $\alpha_{\perp}$ what is suitable for observation, both experimental and computational, does not invalidate a good phenomenological theory for analysis but care should be exercised when attempting to interpret the experimental behavior in terms of ill-defined experimental observables.
Acknowledgements.
We acknowledge the Basque Government for financial support through grants IT1254-19, IT1584-22 and SGi/IZO-SGIker UPV/EHU for computational resources. Funding from the Spanish MCIN/AEI/10.13039/501100011033 is also acknowledged, through grants PID2019-107338RB-C61 and PID2022-139776NB-C65, as well as a María de Maeztu award to Nanogune, Grant CEX2020-001038-M, and the United Kingdom’s EPSRC Grant no. EP/V062654/1. MVFS was funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award No. DE-SC0019394, as part of the CCS Program. FB thanks the ICL RCS High-Performance Computing facility and the UK Materials and Molecular Modelling Hub, partially funded by the EPSRC (Grant Nos. EP/P020194/1 and EP/T022213/1). We thank Dr. Óscar Pozo and Dr. Jon Romero for useful discussions.